# On a discrete version of the position-momentum commutation relation

### v4

@@ XIII Mehta states
$\begin{array}[]{l}\Psi_{n}(q)\!=\!\frac{1}{\sqrt{n!\,2^{n}\sqrt{\pi}}}\,H_{n}(q)\ \mathrm{Exp}[-\frac{1}{2}q^{2}]\end{array}$ (75)
@@
$\hat{F}\Psi_{n}\!=\!(-\mathrm{i})^{n}\Psi_{n}.$ (77)
In his article [7], Mehta has defined the discrete version $\psi_{M,n}\!:\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{R}$,
$\begin{array}[]{l}\psi_{M,n}(m)\!=\!\sum\limits_{\alpha=-\infty}^{\infty}H_{n}(m\!+\!\alpha d)\ \mathrm{Exp}[-\frac{1}{2}(m\!+\!\alpha d)^{2}]\end{array}$ (78)
@@
Table 11: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Mehta states $\psi_{M,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
XIV Harper states
In continuous-variable case, $\hat{p}\!=\!\hat{F}^{\dagger}\hat{q}\hat{F}$, and the harmonic oscillator Hamiltonian can be written as
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{\hat{p}^{2}\!+\!\hat{F}\hat{p}^{2}\hat{F}^{\dagger}}{2}.$ (80)
@@ XIV Harper states
$\begin{array}[]{l}\hat{\mathfrak{F}}D^{2}\hat{\mathfrak{F}}^{\dagger}\psi(n)\!=\!\left(2\cos\frac{2\pi n}{d}\!-\!1\right)\psi(n).\end{array}$ (82)
@@
$\hat{\mathfrak{H}}_{d}\!=\!-\frac{1}{2}(D^{2}\!+\!\hat{\mathfrak{F}}D^{2}\hat{\mathfrak{F}}^{\dagger})$ (83)
@@
and are called Harper functions [8].
For any small $\varepsilon\!>\!0$, some states $\psi_{H,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 12), in the case d=31d\!=\!31, the eigenstates $\psi_{H,0}$, $\psi_{H,1}$, $\psi_{H,2}$ , $\psi_{H,3}$ belong to $\mathcal{S}_{0.001}$.
Table 12: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Harper states $\psi_{H,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
XV Kravchuk states
@@
The Kravchuk functions [9, 10]
@@
Table 13: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Kravchuk states $\psi_{K,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
XVI Concluding remarks
In the case of a continuous-variable quantum system, the canonical position-momentum commutation relation is exactly satisfied in a subspace of the Hilbert space, and plays a fundamental role. The discrete version of this important relation is also satisfied in a subspace, but not exactly, and this happens only if the dimension of the Hilbert space is high enough. We have investigated the behaviour with respect to the discrete version of some important states used in the mathematical description of qudits, including discrete Gaussian states, discrete coherent states and discrete versions of the Hermite-Gauss states. We think that their property discovered in this way is important enough, and may have certain applications.
Quantum Information uses a continuous-variable description, but it is mainly limited to a very particular class of states (including the Gaussian states) and a very particular class of unitary transforms (including the Gaussian unitaries). The discrete-variable version of the commutation relation of the discrete-variable version of position and momentum operator is acceptably satisfied by certain particular states. We think that, some useful models involving only such particular states, can be built in the discrete-variable case by following the analogy with the continuous-variable case. A first attempt in this direction can be found in [4].

### v5

@@ IV.7 Mehta states
$\begin{array}[]{l}\Psi_{n}(q)\!=\!\frac{1}{\sqrt{n!\,2^{n}\sqrt{\pi}}}\,H_{n}(q)\ \mathrm{e}^{-\frac{1}{2}q^{2}}\end{array}$ (94)
@@
$F\Psi_{n}\!=\!(-\mathrm{i})^{n}\Psi_{n}.$ (96)
In his article Mehta87 , Mehta has defined the discrete version $\psi_{M,n}\!:\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{R}$,
$\begin{array}[]{l}\psi_{M,n}(m)\!=\!\sum\limits_{\alpha=-\infty}^{\infty}H_{n}(m\!+\!\alpha d)\ \mathrm{e}^{-\frac{1}{2}(m\!+\!\alpha d)^{2}}\end{array}$ (97)
@@
Table 11: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Mehta states $\psi_{M,k}$.
$\psi$   $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
IV.8 Harper states
In continuous-variable case, $\hat{p}\!=\!F^{\dagger}\hat{q}F$, and the harmonic oscillator Hamiltonian can be written as
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{\hat{p}^{2}\!+\!F\hat{p}^{2}F^{\dagger}}{2}.$ (99)
@@ IV.8 Harper states
$\begin{array}[]{l}\mathbf{F}D^{2}\mathbf{F}^{\dagger}\psi(n)\!=\!\left(2\cos\frac{2\pi n}{d}\!-\!1\right)\psi(n).\end{array}$ (101)
@@
$\hat{\mathbf{H}}_{d}\!=\!-\frac{1}{2}(D^{2}\!+\!\mathbf{F}D^{2}\mathbf{F}^{\dagger})$ (102)
@@
and are called Harper functions Barker .
For any small $\varepsilon\!>\!0$, some states $\psi_{H,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 12), in the case d=31d\!=\!31, the eigenstates $\psi_{H,0}$, $\psi_{H,1}$, $\psi_{H,2}$ , $\psi_{H,3}$ belong to $\mathcal{S}_{0.001}$. Again, there is an almost equidistant distribution in the lower part of the spectrum (see Fig. 6). The Hamiltonian (102) is an other discrete version of (77), and a behavior similar to (76) is again expected to happen.
Table 12: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Harper states $\psi_{H,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
Figure 6: Eigenvalues of the Hamiltonian (102) in the case d=31d\!=\!31.
IV.9 Kravchuk states
@@
The Kravchuk functions Nikiforov ; Vilenkin
@@
Table 13: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain Kravchuk states $\psi_{K,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
V Some possible physical interpretations
We have proved, mainly numerically, that in the case of a quantum system (qudit) described by a $d$-dimensional Hilbert space, a position-like operator $\hat{\mathbf{q}}$ and a momentum-like operator $\hat{\mathbf{q}}$ satisfy $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi$ for certain states $\psi$ if the dimension $d$ is high enough. A fundamental question concerns the existence of any physical interpretation for the operators $\hat{\mathbf{q}}$ and $\hat{\mathbf{p}}$ defined mathematically by following the analogy with the continuous-variable case. It is known that, in the case of a quantum system with $SU(2)$ symmetry, the spectrum of $J_{z}$ has the form $\{-s,-s\!+\!1,..,s\!-\!1,s\}$. So, for example, we can choose $\hat{\mathbf{q}}\!=\!J_{z}$ and $\hat{\mathbf{p}}$ to be the Fourier conjugate operator. More general, any observable with equidistant eigenvalues can play the role of $\hat{\mathbf{q}}$. For eigenvectors forming a basis of the Hilbert space there exists a privileged order, namely the order for which the corresponding eigenvalues form an increasing sequence.
The appearance of quantum states which satisfy with approximation the discrete version of the position-momentum commutation relation among the eigenstates with small energy of certain Hamiltonians suggests a possibility to experimentally obtain such states. They occur naturally when we measure a discrete version of the harmonic quantum oscillator Hamiltonian.
VI Concluding remarks
In the case of a continuous-variable quantum system, the canonical position-momentum commutation relation is exactly satisfied in a subspace of the Hilbert space, and plays a fundamental role. The discrete version of this important relation is also satisfied in a subspace, but not exactly, and this happens only if the dimension of the Hilbert space is high enough. We have investigated the behavior with respect to this discrete version of some important states used in the mathematical description of qudits, including discrete Gaussian states, discrete coherent states and discrete versions of the Hermite-Gauss states. Thus, we have obtained a new property (to our knowledge) of these important quantum states. The discrete version of the position-momentum is satisfied with approximation of certain of them if the dimension of the Hilbert space is high enough. The numerical data presented in this article (Figs. 1-6 and Tables I-XIII) have been obtained using the computer programs presented at the end of the preprint Cotfas-arXiv .
At this moment, there is no satisfactory mathematical explanation for the behavior of the analyzed quantum states concerning the discrete version of the position-momentum commutation relation. Why there exist states for which the relation is satisfied with approximation ? Why they are among the the low energy states ? Why their existence depends on the dimension of the Hilbert space ? There seems to exist a connection between the approximate satisfaction of the position-momentum commutation relation and the tendency of corresponding eigenvalues to be distributed equidistant.
Quantum Information uses a continuous-variable description, but it is mainly limited to a very particular class of states (including the Gaussian states) and a very particular class of unitary transforms (including the Gaussian unitaries). The discrete-variable version of the commutation relation of position -momentum operators is acceptable satisfied by certain particular states. We think that, some useful models involving only such particular states, can be built in the discrete-variable case by following the analogy with the continuous-variable case.
Qudit-based quantum computing Wang2020 is an alternative to the conventional qubit-based quantum computing. The use of quantum systems described by Hilbert spaces of dimension d>2d\!>\!2 provides a larger space to store and process information, and the ability to do multiple control operations simultaneously. These features may allow a reduction of the circuit complexity, a simplification of the experimental setup, and offer the possibility to use more efficient algorithms.